A sample of gas of mass 3.25 g occupies a volume of 426 mL at 0 degree Celcius and 1.75 atm pressure. What is the molecular weight of the gas? ___g/mol

1
Expert's answer
2021-10-30T01:11:02-0400

PV= nRT

1.75×426=n×8.314×273

745.5= 2269.722n


n= 745.5/2269.722

= 0.33

0.33= 3.25

1=. ?

1×3.25/0.33= 9.85g/mol